OpenStudy (anonymous):

-(5s+6)=2(3s=8) what is s?

OpenStudy (anonymous):

is the equation -(5s + 6) = 2 (3s + 8) or is it -(5s + 6) = 2 (3 - 8)??? you put an equal sign in the middle of the 3s and the 8

OpenStudy (anonymous):

Assuming that the equation is -(5s + 6) = 2 (3s + 8) Then here is how you solve it: -(5s + 6) = 2 (3s + 8) -5s - 6 = 6s + 16 -5s - 6 + 6 = 6s + 16 + 6 -5s = 6s + 22 -5s - 6s = 6s - 6s + 22 -11s = 22 -11s / -11 = 22 / -11 s = -2